# On the Role of Attention Masks and LayerNorm in Transformers

### v1

@@ Abstract
Self-attention is the key mechanism of transformers, which are the essential building blocks of modern foundation models. Recent studies have shown that pure self-attention suffers from an increasing degree of rank collapse as depth increases, limiting model expressivity and further utilization of model depth. The existing literature on rank collapse, however, has mostly overlooked other critical components in transformers that may alleviate the rank collapse issue. In this paper, we provide a general analysis of rank collapse under self-attention, taking into account the effects of attention masks and layer normalization (LayerNorm). In particular, we find that although pure masked attention still suffers from exponential collapse to a rank one subspace, local masked attention can provably slow down the collapse rate. In the case of self-attention with LayerNorm, we first show that for certain classes of value matrices, collapse to a rank one subspace still happens exponentially. However, through construction of nontrivial counterexamples, we then establish that with proper choice of value matrices, a general class of sequences may not converge to a rank one subspace, and the self-attention dynamics with LayerNorm can simultaneously possess a rich set of equilibria with any possible rank between one and full. Our result refutes the previous hypothesis that LayerNorm plays no role in the rank collapse of self-attention and suggests that self-attention with LayerNorm constitutes a much more expressive, versatile nonlinear dynamical system than what was originally thought.
@@ 1 Introduction
Transformers are known to exhibit the rank collapse phenomenon [^1] This is sometimes also referred to as the oversmoothing [27, 13, 31] or the token uniformity [10, 22] problem. [10, 22, 13, 27, 31, 17], which refers to the observation that increasing the number of self-attention layers leads to homogeneous token representations. To gain more insight into the rank collapse issue and better understand the effects of self-attention in multi-layer models, it is essential to study the long-term behavior of tokens under self-attention dynamics [10, 13, 14].
However, many existing studies do not take into account architectural components that are commonly used in practice. For instance, first, the theoretical analysis of long-term self-attention dynamics often assumes that attention is fully bidirectional — that is, all tokens are allowed to attend to all other tokens in the sequence [10, 22, 14, 13], which only applies to certain attention mechanisms such as the one deployed in the BERT family [9, 20]. The vast majority of popular transformer architectures used nowadays in practice, including the GPT models [24, 6], use the causal attention masks where tokens are only allowed to attend to preceding tokens, or have sparse attention structure (sparse transformers) [34, 35, 4, 8, 26, 18], where attention is restricted to local interactions between tokens and their chosen neighbors. This limits the practical applicability of the theoretical results developed in [10, 22, 14, 13], as they heavily rely on the key assumption that attention is fully bidirectional.
@@ 3 Problem Setup
Throughout the analysis in the paper, we formalize the attention mask to be a directed graph $\mathcal{G}$. Formally, we represent a directed graph with $N$ nodes by $\mathcal{G}$ and let $E(\mathcal{G})$ be the set of directed edges of $\mathcal{G}$. If there is a directed edge going from $j$ to $i$ in $\mathcal{G}$, for the attention mechanism it means that token $j$ serves as a direct context for token $i$ or token $i$ attends to token $j$, then $(j,i)\in E(\mathcal{G})$. The set $\mathcal{N}_{i}$ of all neighbors of node $i$ is then $\{k|(k,i)\in E(\mathcal{G})\}$.
Furthermore, we will be using the following graph-theoretic terminologies:
@@
Definition 5 (Diameter).
The diameter of a quasi-strongly connected graph is defined to be the longest distance from a center node to any node in the graph. If there are multiple center nodes, then it is the smallest longest distance among them.
@@ 3.2 LayerNorm
Given token representations $X\in\mathbb{R}^{N\times d}$, LayerNorm subtracts the mean across different columns in each row and then scales each row to have a unit $2$-norm. In this work, we consider LayerNorm to only perform the scaling operation, which is a common assumption in theoretical analyses of the attention mechanism [13, 28]. Mathematically, let $D=\operatorname{diag}(d_{1},d_{2},...,d_{d})$ where $d_{i}=1/\|X_{i,:}\|_{2}$ for all $i\in[N]$, then $\operatorname{LN}(X)=DX\,.$
@@ 4.1 Masked Attention
where $r$ is the diameter of $\mathcal{G}$, meaning that tokens converge to a common vector exponentially.
@@
The exponential rate $(1-\epsilon^{r})^{1/r}$ is monotone in the graph diameter $r$. This means that rank collapse should be slower for graphs with larger diameter $r$. Our result thus indirectly supports the use of local attention patterns [35, 4], which not only make the attention computation more efficient (what those works are originally motivated by), but also implicitly alleviate the rank collapse issue.
@@
Finally, for strongly connected graphs, the above result also reveals a trade-off between universal function approximation power and the rate of rank-collapse. Yun et al. [34] show that transformers with strongly connected graph masks are sequence-to-sequence function universal, yet with a mask $\mathcal{G}$ they need at least the diameter of $\mathcal{G}$ layers to achieve the full sequence-to-sequence function approximation property. This implies that masks with smaller diameters are more efficient in terms of function approximation power, yet they are more prone to rank collapse.
@@ 4.2 Masked Attention with LayerNorm: Rank Collapse
Consider the self-attention dynamics with LayerNorm defined in (3). Let $\mathcal{G}$ be a strongly connected graph. Assume A1-A2, and that $W_{V}^{(t)}$ is orthogonal for all $t\geq 0$, and in addition, the initial input $X^{(0)}$ satisfies the following conditions:
  • (i)
    There exists $v\in\mathbb{S}^{d-1}$ such that $\langle X^{(0)}_{i,:},v\rangle>0$ for all $i\in[N]$;
  • (ii)
@@
where $r$ is the diameter of $\mathcal{G}$, meaning that tokens converge to a common point on $\mathbb{S}^{d-1}$ exponentially.
The detailed proof is provided in Appendix C. The result can be seen as a generalized discrete version of Theorem 4.1 in [13]. Notably, our analysis is based purely on advanced linear algebra tools: infinite products of non-negative matrices and their ergodicity, and can account for time-varying weights and general attention masks, as opposed to fixed $W_{K},W_{Q},W_{V}$ over time and $\mathcal{G}$ being complete (which is the case in [13]). One way to satisfy the two conditions on the initial input $X^{(0)}$ is to require $N\leq d$ and to initialize tokens uniformly randomly on $\mathbb{S}^{d-1}$, then the two conditions hold almost surely. This is how the conditions are dealt with in [13].
If the initial token geometry satisfies a stronger condition than (i), then (ii) is no longer necessary and Theorem 2 even directly generalizes to quasi-strongly connected graphs $\mathcal{G}$. We define $:\phi^{(t)}\vcentcolon=\underset{i,j\in[N]}{\min}\langle X^{(t)}_{i,:},X^{(t)}_{j,:}\rangle$, indicating the minimal cosine similarity between tokens. If the cosine similarities are non-negative for all pairs of tokens initially, then the rank collapse happens exponentially, as long as $\mathcal{G}$ is quasi-strongly connected.
@@
$\mu(X^{(t)})\leq C(1-N\epsilon^{2r})^{\frac{t}{2r}},\qquad\forall t\geq 0\,.$ (8)
where $r$ is the diameter of $\mathcal{G}$, meaning that tokens converge to a common point on $\mathbb{S}^{d-1}$ exponentially.
While the requirement in the above result on the initial token geometry might seem restrictive, in practice, such an initial token geometry is indeed satisfied in some commonly used transformer models such as BERT [9] or ALBERT [20] after the initial embedding layer. See the empirical measurements in Section G.4.
@@ 4.4 Discussion
Next, we discuss two further intriguing findings from our analysis of the self-attention dynamics with LayerNorm presented in the previous section.
@@ 5 Numerical Experiments
In this section, we validate our theoretical findings via numerical experiments. Following [10], we randomly select 32 samples of 128-token excerpts (based on the BERT tokenizer) from Wikipedia using the Wikipedia API in Python. More experimental details and additional results can be found in Appendix G.
@@
We first verify the effect of different attention masks and LayerNorm on rank collapse with respect to $\mu(\cdot)$. We use BERT [9] as the backbone transformer model and consider three different model variants for a controlled experiment: self-attention network (SAN) in which the model has self-attention layers, SAN with LayerNorm where in each layer there is self-attention followed by LayerNorm, and finally the standard transformer with all the components. For attention masks, we select four representative types: the complete graph, the causal graph, the sliding window (tokens can only attend to the token right before and after and themselves) and the uni-directional sliding window (tokens can only attend to the token right before and themselves). Fig. 2 shows the average values of $\mu(X^{(t)})$ with standard deviations over the $32$ samples in different architectures. We see that as our theory predicts, in SANs, $\mu(X^{(t)})$ converges to zero exponentially for all attention masks, yet more local attention masks slow down the convergence rate. Furthermore, right after we add in LayerNorm, $\mu(X^{(t)})$ no longer converges to zero and can be maintained significantly above zero, which holds for both initialized and pretrained models. Results in $128$-layer initialized models also exhibit the same trend, and can be found in Section G.1.
Figure 2: Evolution of $\mu(X^{(t)})$ (in log-log scale) as the number of layers increases. Rank collapse happens exponentially for pure attention, despite different attention masks having different convergence rates, as our theory predicts. However, as soon as we solely add in LayerNorm, $\mu(X^{(t)})$ no longer converge to zero in both randomly initialized and pretrained models.
@@
Next, we investigate the effect of the temperature term $d_{QK}$ in the attention calculation. Since the amount of focus of the attention is affected by both the attention mask and the temperature term, we investigate the effect of $d_{QK}$ with different attention masks. Fig. 3 shows the average values of $\mu(X^{(t)})$ with standard deviations over the $32$ samples with different masks in pretrained SANs. We observe that while a smaller temperature term $d_{QK}$ alleviates the (initial) rate of rank collapse in all cases, the effect is more significant with global attention than with sparser masked attention, and more in shallower layers than in deeper layers. The results in initialized models show similar trends and can be found in Section G.2.
@@
Finally, we study the evolution of token geometry as the number of layers increases in transformers. Specifically, to capture more fine-grained details than $\mu(\cdot)$, we measure how the rank, minimal singular value of token representations, and absolute values of pairwise cosine similarities among tokens change as tokens pass through the transformer layers. We select four representative pretrained transformer models: BERT [9], GPT2 [24], T5 [25] and ALBERT [20] and the average results over $32$ samples are shown in Fig. 4. We see that all models exhibit similar behaviors as tokens evolve along the layers: the models can effectively preserve the full rank of token representations, as also evident from the minimal singular values consistently remaining significantly above zero. Yet the absolute cosine similarities among tokens suggest that tokens gradually concentrate in a narrow region. This empirical observation aligns with the characteristics of the equilibrium in our counterexample that a stable long-term geometry for token representations in transformers can retain full rank while being close to a low dimensional geometry at the same time.
Figure 4: Evolution of token geometry as the number of layers increases. We see that tokens are indeed able to maintain full rank, while at the same time the representations are anisotropic, meaning that they concentrate in a narrow region, as indicated by the average pairwise absolute cosine similarities. This is in agreement with our theoretical results about LayerNorm.

### v2

@@ Abstract
Self-attention is the key mechanism of transformers, which are the essential building blocks of modern foundation models. Recent studies have shown that pure self-attention suffers from an increasing degree of rank collapse as depth increases, limiting model expressivity and further utilization of model depth. The existing literature on rank collapse, however, has mostly overlooked other critical components in transformers that may alleviate the rank collapse issue. In this paper, we provide a general analysis of rank collapse under self-attention, taking into account the effects of attention masks and layer normalization (LayerNorm). In particular, we find that although pure masked attention still suffers from exponential collapse to a rank one subspace, sparse or local masked attention can provably slow down the collapse rate. In the case of self-attention with LayerNorm, we first show that for certain classes of value matrices, collapse to a rank one subspace still happens exponentially. However, through construction of nontrivial counterexamples, we then establish that with proper choice of value matrices, a general class of sequences may not converge to a rank one subspace, and the self-attention dynamics with LayerNorm can simultaneously possess a rich set of equilibria with any possible rank between one and full. Our result refutes the previous hypothesis that LayerNorm plays no role in the rank collapse of self-attention and suggests that self-attention with LayerNorm constitutes a much more expressive, versatile nonlinear dynamical system than what was originally thought.
@@ 1 Introduction
Transformers are known to exhibit the rank collapse phenomenon [^1] This is also referred to as the oversmoothing [33, 15, 37] or the token uniformity [12, 27] problem. [12, 27, 15, 33, 37, 20], which refers to the observation that increasing the number of self-attention layers leads to homogeneous token representations. To gain more insight into the rank collapse issue and better understand the effects of self-attention in multi-layer models, it is essential to study the long-term behavior of tokens under self-attention dynamics [12, 15, 16].
However, many existing studies do not take into account architectural components that are commonly used in practice. For instance, first, the theoretical analysis of long-term self-attention dynamics often assumes that attention is fully bidirectional — that is, all tokens are allowed to attend to all other tokens in the sequence [12, 27, 16, 15], which only applies to certain attention mechanisms such as the one deployed in the BERT family [11, 24]. The vast majority of popular transformer architectures used nowadays in language models, including the GPT models [30, 6], use the causal attention masks where tokens are only allowed to attend to preceding tokens, or have sparse attention structure (sparse attention) [41, 42, 4, 10, 32, 21, 38, 22, 8], where attention is restricted to local interactions between tokens and their chosen neighbors. This limits the practical applicability of the theoretical results developed in [12, 27, 16, 15], as they heavily rely on the key assumption that attention is fully bidirectional.
@@ 2 Related work
Sparse and local Attention
While many existing transformer models and LLMs do not use sparse attention, sparse and local attention is gaining popularity due to the demand for efficiency, particular for long-context tasks. For example, sparse attention was populated by Longformer [4] and OpenAI [10] and nowadays popular LLMs like Mistral 7B [22] use sliding window attention by default. Other popular sparse attention models include, but are not limited to BigBird [42], Recurrent Memory Transformers (RMTs) [7], and Streaming Attention [38]. Besides language tasks, sparse attention is also common in vision transformers [25, 28, 19].
@@ 3 Problem Setup
Throughout the analysis in the paper, we formalize the attention mask to be a directed graph $\mathcal{G}$. Formally, we represent a directed graph with $N$ nodes by $\mathcal{G}$ and let $E(\mathcal{G})$ be the set of directed edges of $\mathcal{G}$. If there is a directed edge going from $j$ to $i$ in $\mathcal{G}$, i.e. $(j,i)\in E(\mathcal{G})$, for the attention mechanism it means that token $j$ serves as a direct context for token $i$ or token $i$ attends to token $j$. The set $\mathcal{N}_{i}$ of all neighbors of node $i$ is then $\{k:(k,i)\in E(\mathcal{G})\}$.
Furthermore, we will be using the following graph-theoretic terminology:
@@
Definition 5 (Radius).
The radius of a quasi-strongly connected graph is defined to be the longest distance from a center node to any node in the graph. If there are multiple center nodes, then it is the smallest value among them.
@@ 3.2 LayerNorm
Given token representations $X\in\mathbb{R}^{N\times d}$, LayerNorm subtracts the mean across different columns in each row and then scales each row to have a unit $2$-norm. In this work, we consider LayerNorm to only perform the scaling operation, which is a common assumption in theoretical analyses of the attention mechanism [15, 34] [^2] This definition of LayerNorm, precisely concides with the RMSNorm [43] that is widely deployed in mainstream LLMs. We choose the terminology following the convention of literature.. Mathematically, let $D=\operatorname{diag}(d_{1},d_{2},...,d_{N})$ where $d_{i}=1/\|X_{i,:}\|_{2}$ for all $i\in[N]$, then $\operatorname{LN}(X)=DX\,.$
@@ 4.1 Masked Attention
where $r$ is the radius of $\mathcal{G}$, meaning that tokens converge to a common vector exponentially.
@@
The exponential rate $(1-\epsilon^{r})^{1/r}$ is monotone in the graph radius $r$. This means that rank collapse should be slower for graphs with larger radius $r$. Our result thus indirectly supports the use of local attention patterns [42, 4], which not only make the attention computation more efficient (what those works are originally motivated by), but also implicitly alleviate the rank collapse issue.
@@
Finally, for strongly connected graphs, the above result also reveals a trade-off between universal function approximation power and the rate of rank-collapse. Yun et al. [41] show that transformers with strongly connected graph masks are sequence-to-sequence function universal, yet with a mask $\mathcal{G}$ they need at least the diameter of $\mathcal{G}$ layers to achieve the full sequence-to-sequence function approximation property. This implies that masks with smaller diameters (and thus smaller radii, as radius $\leq$ diameter $\leq$ 2 radius) are more efficient in terms of function approximation power, yet they are more prone to rank collapse.
@@ 4.2 Masked Attention with LayerNorm: Rank Collapse
Consider the self-attention dynamics with LayerNorm defined in (3). Let $\mathcal{G}$ be a strongly connected graph. Assume A1-A2, and that $W_{V}^{(t)}$ is orthogonal for all $t\geq 0$, and in addition, the initial input $X^{(0)}$ satisfies that
  • ()($\ast$)( ∗ )
@@
where $r$ is the radius of $\mathcal{G}$, meaning that tokens converge to a common point on $\mathbb{S}^{d-1}$ exponentially.
The detailed proof is provided in Appendix C. The result can be seen as a generalized discrete version of Theorem 4.1 in [15]. Notably, our analysis is based purely on advanced linear algebra tools: infinite products of non-negative matrices and their ergodicity, and can account for time-varying weights and general attention masks, as opposed to fixed $W_{K},W_{Q},W_{V}$ over time and $\mathcal{G}$ being complete (which is the case in [15]). One way to satisfy the condition ($\ast$) on the initial input $X^{(0)}$ is to require $N\leq d$ and to initialize tokens uniformly randomly on $\mathbb{S}^{d-1}$, then the condition hold almost surely. This is how the condition is dealt with in [15].
Note that the condition ()($\ast$)( ∗ ) implies that there exists $v\in\mathbb{S}^{d-1}$ such that $\langle X^{(0)}_{i,:},v\rangle>0$ for all $i\in[N]$ by either the hyperplane separation theorem or Farkas’ lemma (see Lemma 6 in Appendix C). If the initial token geometry satisfies a stronger condition than the above, then ($\ast$) is no longer necessary and Theorem 2 even directly generalizes to quasi-strongly connected graphs $\mathcal{G}$. We define $:\phi^{(t)}\vcentcolon=\underset{i,j\in[N]}{\min}\langle X^{(t)}_{i,:},X^{(t)}_{j,:}\rangle$, indicating the minimal cosine similarity between tokens. If the cosine similarities are non-negative for all pairs of tokens initially, then the rank collapse happens exponentially, as long as $\mathcal{G}$ is quasi-strongly connected.
@@
$\mu(X^{(t)})\leq C(1-\epsilon^{2r})^{\frac{t}{2r}},\qquad\forall t\geq 0\,.$ (8)
where $r$ is the radius of $\mathcal{G}$, meaning that tokens converge to a common point on $\mathbb{S}^{d-1}$ exponentially.
Full mask vs. causal mask
We can refine Corollary 1 by specifying the number of center nodes $n$ in the mask $\mathcal{G}$, then the upper bound for the exponential rate would be $(1-n\epsilon^{2r})$ instead, meaning that the rate of rank collapse can be negatively affected by the number of center nodes in the mask $\mathcal{G}$. In the case of full attention where $\mathcal{G}$ is the complete graph, the mask would have $N$ center nodes, matching the bound in Theorem 2. In the case of causal attention where $\mathcal{G}$ is the causal graph, the mask would only have one center node, and the upper bound would be looser, suggesting the advantage of the causal mask in mitigating the rate of rank collapse as compared to the full mask.
Post-LN vs. Pre-LN
The definition of LayerNorm in (3) follows from the original transformer paper [35] and nowadays it is referred as post-LN [39]. An alternative use of LayerNorm in many LLMs, where LayerNorm comes before self-attention, is called pre-LN and can be written instead as
$X^{(t+1)}=A^{(t)}\operatorname{LN}({X}^{(t)})W^{(t)}\vcentcolon=A^{(t)}D^{(t)}X^{(t)}W_{V}^{(t)}\,.$ (9)
Note that Theorem 2 and Corollary 1 apply directly to the case of pre-LN with similar proofs.
@@ 4.3.1 Illustrative Counterexample
Remark 2.
For any orthogonal $Z\in\mathbb{R}^{d\times d}$, $W_{Z}=Z^{\top}WZ$ works equivalently in this example, and the resulting token trajectories follow $X^{(t)}Z$.
@@ 4.4 Discussion
Next, we discuss three further intriguing findings from our analysis of the self-attention dynamics with LayerNorm presented in the previous section.
Scaling effect of LayerNorm is key
From a dynamical system perspective, rank collapse happens under pure self-attention because the attraction of the center node causes all the tokens to be aligned. In the counterexample provided in Section 4.3.1, the key insight to prevent the second token from aligning with the first token is that the updated representation of the second token must generate a component canceling out the attraction from the first token, which acts as a repulsion. The crucial role of LayerNorm here is to stabilize this cancellation process by readjusting the scale of tokens to ensure that the cancellation persists in all the updates, which would not be the case if there is no LayerNorm.
@@ 5 Numerical Experiments
In this section, we validate our theoretical findings via numerical experiments. Following [12], we randomly select 3000 samples of 128-token excerpts (based on the BERT tokenizer) from Wikipedia using the Wikipedia API in Python. More experimental details and additional results can be found in Appendix G.
@@
We first verify the effect of different attention masks and LayerNorm on rank collapse with respect to $\mu(\cdot)$. We use BERT [11] as the backbone transformer model and consider five different model variants for a controlled experiment: self-attention network (SAN) in which the model has self-attention layers; SAN with skip connections; SAN with LayerNorm where in each layer there is self-attention followed by LayerNorm; SAN with both skip connections and LayerNorm where LayerNorm is put after skip connections; and finally the standard transformer with all the components. For attention masks, we select four representative types: the complete graph, the causal graph, the sliding window (tokens can only attend to the token right before and after and themselves) and the uni-directional sliding window (tokens can only attend to the token right before and themselves). Fig. 2 shows the average values of $\mu(X^{(t)})$ with standard deviations over the $3000$ samples in different architectures. We see that in SANs, $\mu(X^{(t)})$ converges to zero exponentially for all attention masks, yet more local attention masks slow down the convergence rate. Furthermore, in randomly initialized models, right after we add in LayerNorm, $\mu(X^{(t)})$ no longer converges to zero and can be maintained at a stable level significantly above zero. In pretrained models, LayerNorm helps prevent the issue together with other components such as skip connections and stabilize the representations. Results in $128$-layer initialized models also exhibit the same trend, and can be found in Section G.1.
Figure 2: Evolution of $\mu(X^{(t)})$ (in log-log scale) as the number of layers increases. Rank collapse happens exponentially for pure attention, despite different attention masks having different convergence rates. However, as soon as we solely add in LayerNorm, $\mu(X^{(t)})$ no longer converge to zero in randomly initialized models; in pretrained models, LayerNorm helps prevent the issue together with other components and stabilize the representations.
@@
The complex interplay between different components in transformers
Previous works have shown that skip connections can help combat rank collapse in transformers. Yet in Fig. 2, skip connections seem to make $\mu(X)$ unstable, particularly in deeper layers (for 128-layer results, see Section G.1). Compared with full transformers where $\mu(X)$ stays relatively stable, there is a clear discrepancy. In that case, LayerNorm emerges as a crucial element in mitigating rank collapse and stabilizing token representations while also counteracting potential negative effects of pure skip connections. This underscores the complex interplay between different components in transformers.
@@
Next, we investigate the effect of the temperature term $d_{QK}$ in the attention calculation. Since the amount of focus of the attention is affected by both the attention mask and the temperature term, we investigate the effect of $d_{QK}$ with different attention masks. Fig. 3 shows the average values of $\mu(X^{(t)})$ with standard deviations over the $3000$ samples with different masks in pretrained SANs. We observe that while a smaller temperature term $d_{QK}$ alleviates the (initial) rate of rank collapse in all cases, the effect is more significant with global attention than with sparser masked attention, and more in shallower layers than in deeper layers. The results in initialized models show similar trends and can be found in Section G.2.
@@
Finally, we study the evolution of token geometry as the number of layers increases in transformers. Specifically, to capture more fine-grained details than $\mu(\cdot)$, we measure how the rank, minimal singular value of token representations, and absolute values of pairwise cosine similarities among tokens change as tokens pass through the transformer layers. We select four representative pretrained transformer models: BERT [11], GPT2 [30], T5 [31] and ALBERT [24] and the average results over $3000$ samples are shown in Fig. 4. We see that all models exhibit similar behaviors as tokens evolve along the layers: the models can effectively preserve the full rank of token representations, as also evident from the minimal singular values consistently remaining significantly above zero. Yet the absolute cosine similarities among tokens suggest that tokens gradually concentrate in a narrow region. This empirical observation aligns with the characteristics of the equilibrium in our counterexample that a stable long-term geometry for token representations in transformers can retain full rank while being close to a low dimensional geometry at the same time.
Figure 4: Evolution of token geometry as the number of layers increases. We see that tokens are indeed able to maintain full rank, while at the same time the representations are anisotropic, meaning that they concentrate in a narrow region, as indicated by the average pairwise absolute cosine similarities.